OpenStudy (anonymous):

The length of a box is 5 more than the width and the height is 2 less than the width. If the volume is equal to 150 cubic feet, what are the measurements for each side.

OpenStudy (anonymous):

@dtan5457 @mathrulezz @aum @Abmon98 @pooja195 @welshfella @ineptAtMath

OpenStudy (welshfella):

let the width be x , then height = x -2 and the length = x + 5 vilume = w*h*l so x(x-2)(x + 5) = 150 you need to solve this for x

OpenStudy (anonymous):

I dont think i did it right i ot x^3+3x^2-10x=150

OpenStudy (welshfella):

x^3 + 3x^2 - 10x - 150 = 0

OpenStudy (welshfella):

these are not straight forward to solve but as the last term is 150 and 5 is a factor try plugging x = 5 into the equation and see if you get 0

OpenStudy (welshfella):

125 =75 - 50 - 100 = 0!!!!

OpenStudy (anonymous):

yea that works

OpenStudy (welshfella):

* 125 + 75

OpenStudy (welshfella):

got it1 so width = 5 and you can work out the rest easily

OpenStudy (anonymous):

so the length is 10 right?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

great thanks

OpenStudy (welshfella):

yw